# Odd-Parity Magnetism and Gate-Tunable Edelstein Response in van der Waals Heterostructures

### v1

@@ Abstract
where $n_{\textrm{F}}$ and $\partial_{\epsilon}n_{\textrm{F}}$ denotes the Fermi-Dirac distribution function and its derivative respectively. The explicit forms of the prefactors appearing in Eq. (4) are given by $A_{\mathbf{k}}=\frac{\lambda^{4}\varepsilon_{\mathbf{k+Q}}^{2}}{16\Delta_{\mathbf{k}}^{2}}$ and $B_{\mathbf{k}}=\frac{\lambda^{4}\varepsilon_{\mathbf{k+Q}}^{2}}{16\varepsilon_{\mathbf{k}}\Delta_{\mathbf{k}}^{2}}+\frac{\lambda^{4}\varepsilon_{\mathbf{k}}\varepsilon_{\mathbf{k+Q}}^{2}}{8\Delta_{\mathbf{k}}^{3}}$, with $\Delta_{\mathbf{k}}\equiv\varepsilon_{\mathbf{k}}^{2}-\varepsilon_{\mathbf{k+Q}}^{2}$. (See Sec. II of the Supplemental Material for the detailed derivation.) The Fermi-surface term restricts the contribution to the near Fermi surface, whereas the Fermi-sea term integrates over the entire BZ. This decomposition highlights a competition between two distinct instability mechanisms.
@@ 1 Analytical calculation of coefficient $\beta_{2}$
$B^{(\mathbf{k})}(\mathbf{k})=\frac{\lambda^{4}\varepsilon_{\mathbf{k+Q}}^{2}}{16\varepsilon_{\mathbf{k}}\Delta_{\mathbf{k}}^{2}}+\frac{\lambda^{4}\varepsilon_{\mathbf{k}}\varepsilon_{\mathbf{k+Q}}^{2}}{8\Delta_{\mathbf{k}}^{3}},\quad\text{}B^{(\mathbf{k+Q})}(\mathbf{k})=\frac{\lambda^{4}\varepsilon_{k}^{2}}{16\varepsilon_{\mathbf{k+Q}}\Delta_{\mathbf{k}}^{2}}-\frac{\lambda^{4}\varepsilon_{\mathbf{k+Q}}\varepsilon_{k}^{2}}{8\Delta_{\mathbf{k}}^{3}}$ (S31)
@@ IV Symmetry of the tight-binding model
The commensurate stripe order is characterized by the wavevector $\mathbf{Q}=(\pi,0)$, which breaks the primitive translational symmetry and doubles the unit cell. As a result, the Bloch Hamiltonian acts on an $8$-dimensional internal space ($2\times 2\times 2$), arising from the sublattice, band-folding (due to unit-cell doubling), and spin degrees of freedom.

### v2

@@ Abstract
where $n_{\textrm{F}}$ and $\partial_{\epsilon}n_{\textrm{F}}$ denotes the Fermi-Dirac distribution function and its derivative respectively. The explicit forms of the prefactors appearing in Eq. (4) are given by $A_{\mathbf{k}}=\frac{\lambda^{4}\varepsilon_{\mathbf{k+Q}}^{2}}{16\Delta_{\mathbf{k}}^{2}}$ and $B_{\mathbf{k}}=\frac{\lambda^{4}\varepsilon_{\mathbf{k+Q}}^{2}}{16\varepsilon_{\mathbf{k}}\Delta_{\mathbf{k}}^{2}}+\frac{\lambda^{4}\varepsilon_{\mathbf{k}}\varepsilon_{\mathbf{k+Q}}^{2}}{4\Delta_{\mathbf{k}}^{3}}$, with $\Delta_{\mathbf{k}}\equiv\varepsilon_{\mathbf{k}}^{2}-\varepsilon_{\mathbf{k+Q}}^{2}$. (See Sec. II of the Supplemental Material for the detailed derivation.) The Fermi-surface term restricts the contribution to the near Fermi surface, whereas the Fermi-sea term integrates over the entire BZ. This decomposition highlights a competition between two distinct instability mechanisms.
@@ 1 Analytical calculation of coefficient $\beta_{2}$
$B^{(\mathbf{k})}(\mathbf{k})=\frac{\lambda^{4}\varepsilon_{\mathbf{k+Q}}^{2}}{16\varepsilon_{\mathbf{k}}\Delta_{\mathbf{k}}^{2}}+\frac{\lambda^{4}\varepsilon_{\mathbf{k}}\varepsilon_{\mathbf{k+Q}}^{2}}{4\Delta_{\mathbf{k}}^{3}},\quad\text{}B^{(\mathbf{k+Q})}(\mathbf{k})=\frac{\lambda^{4}\varepsilon_{k}^{2}}{16\varepsilon_{\mathbf{k+Q}}\Delta_{\mathbf{k}}^{2}}-\frac{\lambda^{4}\varepsilon_{\mathbf{k+Q}}\varepsilon_{k}^{2}}{4\Delta_{\mathbf{k}}^{3}}$ (S31)
@@ IV Symmetry of the tight-binding model
The commensurate stripe order is characterized by the wavevector $\mathbf{Q}=(\pi,0)$, which breaks the primitive translational symmetry and doubles the unit cell. As a result, the Bloch Hamiltonian acts on an 88-dimensional internal space ($2\times 2\times 2$), arising from the sublattice, band-folding (due to unit-cell doubling), and spin degrees of freedom.